# The LHC has ruled out Supersymmetry – really?

### v1

@@ 4 Phenomenological MSSM
When requiring $|\mu|<400$ GeV, as preferred by the naturalness argument elaborated on in the previous section, only the region with $m_{\tilde{\chi}^{\pm}_{1}}<400$ GeV remains from Fig. 7, with the density of points considerably thinned out, see Fig. 8.777The low density of points with $|\mu|<400$ GeV is simply a volume effect from the sampling of a vast parameter space. The overall picture, however, remains unchanged, with many well-motivated light SUSY scenarios escaping current limits. Specifically one finds along the diagonal many higgsino LSP points with very low $r^{\rm comb}_{\rm obs}$, which may be challenging to detect also in future LHC runs; these typically have heavy winos ($|M_{2}|>1$ TeV). The higgsino LSP points with a promising $r^{\rm comb}_{\rm obs}$ are typically accompanied by winos below about 1 TeV. There are also points with wino-like LSPs, but with a large enough higgsino or bino admixture to make the $\tilde{\chi}^{\pm}_{1}$ not too long-lived (or even decay promptly). As pointed out in Carpenter et al. (2024), part of this parameter space could be tested by hadronic mono-boson searches, in particular if the MET cuts are optimised. Away from the diagonal, that is for sizeable $\tilde{\chi}^{\pm}_{1}-\tilde{\chi}^{0}_{1}$ mass differences, the spectrum consists of light higgsinos with a bino LSP; again the $r$-values are higher when winos are not too heavy, so they also contribute to the signal. Details will be discussed in Constantin et al. (2025).
@@ 5 Hints for light electroweak-inos?
Last but not least, we note that there are tantalising, small but seemingly consistent, excesses in ATLAS and CMS supersymmetry searches involving two or more soft leptons, see Aad et al. (2020, 2021a); Tumasyan et al. (2022); Hayrapetyan et al. (2024). The excesses mostly appear in the dilepton ($2\ell$+MET) channel, with a dilepton invariant mass in the range of $10-20$ GeV, and can be interpreted as stemming from light charginos/neutralinos, $pp\to\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{2}\to W^{*}\tilde{\chi}^{0}_{1}Z^{*}\tilde{\chi}^{0}_{1}$, with a compressed mass spectrum, roughly around $\Delta m=m_{\tilde{\chi}^{0}_{2}}-m_{\tilde{\chi}^{0}_{1}}\approx 10-30$ GeV. Small excesses are also seen in monojet searches [Aad et al. (2021b); Tumasyan et al. (2021)] and, as pointed out in Agin et al. (2024a), may be compatible with the soft lepton ones.
The simplified models used by ATLAS and CMS for their interpretations, a wino/bino model with $m_{\tilde{\chi}^{0}_{2}}=m_{\tilde{\chi}^{\pm}_{1}}$ and a higgsino LSP model with $m_{\tilde{\chi}^{\pm}_{1}}=0.5(m_{\tilde{\chi}^{0}_{2}}+m_{\tilde{\chi}^{0}_{1}})$, are, however, never exactly realised in the MSSM. In the wino/bino case, for the small mass splittings relevant here, there will always be some wino-bino mixing. Moreover, as pointed out by Baum et al. (2023); Chakraborti et al. (2024), $\tilde{\chi}^{0}_{2}\to\tilde{\chi}^{0}_{1}\gamma$ decays may be relevant and weaken the leptonic signal. In the higgsino LSP case, on the other hand, a non-negligible gaugino admixture is needed to make the mass splitting large enough, as pure higgsinos have loop-suppressed mass splittings below about 1 GeV.
Realistic MSSM scenarios for explaining the observed excesses, in part including also dark matter (DM) constraints, were considered in Agin et al. (2024a); Chakraborti et al. (2024); Agin et al. (2024b); Martin (2024). Chakraborti et al. (2024) concluded that wino/bino DM in the MSSM performs better than the higgsino scenario (for $\mu>0$) and for $m_{\tilde{\chi}^{0}_{2}}\approx m_{\tilde{\chi}^{\pm}_{1}}\leq 250$ GeV and $\Delta m=m_{\tilde{\chi}^{0}_{2}}-m_{\tilde{\chi}^{0}_{1}}\approx 25$ GeV, has roughly the right cross section to produce the observed excesses. Martin (2024) considered either sign of $\mu$ and concluded that $\mu<0$ and low $\tan\beta$ (with decoupled heavy Higgs bosons) allows one to evade DM direct detection limits, rendering light higgsinos a viable option. Agin et al. (2024b) discussed the complementarity of the soft lepton excesses with other excesses in LHC monojet searches for the MSSM and next-to-MSSM, as well as for two non-SUSY scenarios (a vector-like lepton model and a type-II see-saw model); they also identified the wino/bino scenario in the MSSM as “the best option currently on the table”, while the non-SUSY interpretations fit the excesses less well.
@@ 6 Summary and conclusions
On the other hand, as discussed in, e.g., Dumont et al. (2014); Arbey et al. (2016, 2022), precise measurements of the Higgs boson offer a valuable and complementary probe of supersymmetry. The presence of additional SUSY particles can modify the couplings of the SM-like Higgs, thereby altering its production cross sections and decay branching ratios. As a result, precision Higgs measurements can impose meaningful constraints on the SUSY parameter space, even in scenarios where direct searches are not effective.

### v2

@@ 4 Phenomenological MSSM
When requiring $|\mu|<400$ GeV, as preferred by the naturalness argument elaborated on in the previous section, only the region with $m_{\tilde{\chi}^{\pm}_{1}}<400$ GeV remains from Fig. 7, with the density of points considerably thinned out, see Fig. 8.101010The low density of points with $|\mu|<400$ GeV is simply a volume effect from the sampling of a vast parameter space. The overall picture, however, remains unchanged, with many well-motivated light SUSY scenarios escaping current limits. Specifically one finds along the diagonal many higgsino LSP points with very low $r^{\rm comb}_{\rm obs}$, which may be challenging to detect also in future LHC runs; these typically have heavy winos ($|M_{2}|>1$ TeV). The higgsino LSP points with a promising $r^{\rm comb}_{\rm obs}$ are typically accompanied by winos below about 1 TeV. There are also points with wino-like LSPs, but with a large enough higgsino or bino admixture to make the $\tilde{\chi}^{\pm}_{1}$ not too long-lived (or even decay promptly). As pointed out in [41], part of this parameter space could be tested by hadronic mono-boson searches, in particular if the MET cuts are optimised. Away from the diagonal, that is for sizeable $\tilde{\chi}^{\pm}_{1}-\tilde{\chi}^{0}_{1}$ mass differences, the spectrum consists of light higgsinos with a bino LSP; again the $r$-values are higher when winos are not too heavy, so they also contribute to the signal. Details will be discussed in [42].
@@ 5 Hints for light electroweak-inos?
Last but not least, we note that there are tantalising, small but seemingly consistent, excesses in ATLAS and CMS supersymmetry searches involving two or more soft leptons [43, 44, 45, 46, 47]. The excesses mostly appear in the dilepton ($2\ell$+MET) channel, with a dilepton invariant mass in the range of about $10-20$ GeV, and can be interpreted, e.g., as stemming from light charginos/neutralinos, $pp\to\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{2}\to W^{*}\tilde{\chi}^{0}_{1}Z^{*}\tilde{\chi}^{0}_{1}$, with a compressed mass spectrum, roughly around $\Delta m=m_{\tilde{\chi}^{0}_{2}}-m_{\tilde{\chi}^{0}_{1}}\approx 10-30$ GeV. Small excesses are also seen in monojet searches [48, 49] and, as pointed out in [50], may be compatible with the soft lepton ones.
The simplified models used by ATLAS and CMS for their interpretations, a wino/bino model with $m_{\tilde{\chi}^{0}_{2}}=m_{\tilde{\chi}^{\pm}_{1}}$ and a higgsino LSP model with $m_{\tilde{\chi}^{\pm}_{1}}=0.5(m_{\tilde{\chi}^{0}_{2}}+m_{\tilde{\chi}^{0}_{1}})$, are, however, never exactly realised in the MSSM. In the wino/bino case, for the small mass splittings relevant here, there will always be some wino-bino mixing. Moreover, as pointed out in [51, 52], $\tilde{\chi}^{0}_{2}\to\tilde{\chi}^{0}_{1}\gamma$ decays may be relevant and weaken the leptonic signal. In the higgsino LSP case, on the other hand, a non-negligible gaugino admixture is needed to make the mass splitting large enough, as pure higgsinos have loop-suppressed mass splittings below about 1 GeV.
Realistic MSSM scenarios for explaining the observed excesses, in part including also dark matter (DM) constraints, were considered in [50, 52, 53, 54]. Reference [52] concluded that wino/bino DM in the MSSM performs better than the higgsino scenario (for $\mu>0$) and for $m_{\tilde{\chi}^{0}_{2}}\approx m_{\tilde{\chi}^{\pm}_{1}}\leq 250$ GeV and $\Delta m=m_{\tilde{\chi}^{0}_{2}}-m_{\tilde{\chi}^{0}_{1}}\approx 25$ GeV, has roughly the right cross section to produce the observed excesses. Reference [54] considered either sign of $\mu$ and concluded that $\mu<0$ and low $\tan\beta$ (with decoupled heavy Higgs bosons) allows one to evade DM direct detection limits, rendering light higgsinos a viable option. In [53] the complementarity of the soft lepton excesses with other excesses in LHC monojet searches was discussed for the MSSM and next-to-MSSM, as well as for two non-SUSY scenarios (a vector-like lepton model and a type-II see-saw model); they also identified the wino/bino scenario in the MSSM as “the best option currently on the table”, while the non-SUSY interpretations fit the excesses less well.
@@ 6 Summary and conclusions
On the other hand, as discussed in [55, 56, 26], precise measurements of the Higgs boson offer a valuable and complementary probe of supersymmetry. The presence of additional SUSY particles can modify the couplings of the SM-like Higgs, thereby altering its production cross sections and decay branching ratios. As a result, precision Higgs measurements can impose meaningful constraints on the SUSY parameter space, even in scenarios where direct searches are not effective.